A packaging firm relies on the production function Q(L,K) = KL + L. Assuming the firm’s optimal input combination is interior (i.e. it uses positive amounts of both inputs), what is its long-run marginal cost function?

a. ???? = 2√???

b. ???? = √ ?? /?

c. ???? = 2√??? – r

d. ???? = (2?/ √??) − r
Answer #1

For optimum input combination:

MPL/w = MPK/r

(K + 1)/w = L/r

Q = wL2/r  

L = (rQ/w)0.5

Q = (K + 1)2 x r/w

K = (wQ/r)0.5 - 1

Total cost = wL + rK

TC = (wrQ)0.5 + (wrQ)0.5 - r

MC = d(TC)/dQ = (wr/Q)0.5

option b is correct